OpenStudy (anonymous):

Simplify the following expression: (x + 6y) - (3x − 10y). If the final answer is written in the form Ax + By, what is the value of A?

OpenStudy (carson889):

x - 3x = -2x 6y - (-10y) = 16y Thus final answer = -2x + 16y, Where A = -2